# exampleloeList of Examples

### v1

@@ Abstract.
Designing preference elicitation (PE) methodologies that can quickly ascertain a user’s top item preferences in a cold-start setting is a key challenge for building effective and personalized conversational recommendation (ConvRec) systems. While large language models (LLMs) constitute a novel technology that enables fully natural language (NL) PE dialogues, we hypothesize that monolithic LLM NL-PE approaches lack the multi-turn, decision-theoretic reasoning required to effectively balance the NL exploration and exploitation of user preferences towards an arbitrary item set. In contrast, traditional Bayesian optimization PE methods define theoretically optimal PE strategies, but fail to use NL item descriptions or generate NL queries, unrealistically assuming users can express preferences with direct item ratings and comparisons. To overcome the limitations of both approaches, we formulate NL-PE in a Bayesian Optimization (BO) framework that seeks to actively elicit natural language feedback to reduce uncertainty over item utilities to identify the best recommendation. Key challenges in generalizing BO to deal with natural language feedback include determining (a) how to leverage LLMs to model the likelihood of NL preference feedback as a function of item utilities and (b) how to design an acquisition function that works for language-based BO that can elicit in the infinite space of language. We demonstrate our framework in a novel NL-PE algorithm, PEBOL, which uses Natural Language Inference (NLI) between user preference utterances and NL item descriptions to maintain preference beliefs and BO strategies such as Thompson Sampling (TS) and Upper Confidence Bound (UCB) to guide LLM query generation. We numerically evaluate our methods in controlled experiments, finding that PEBOL achieves up to 131% improvement in MAP@10 after 10 turns of cold start NL-PE dialogue compared to monolithic GPT-3.5, despite relying on a much smaller 400M parameter NLI model for preference inference.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Personalizationccs: Information systems Language models
@@ 1. Introduction
Personalized conversational recommendation (ConvRec) systems require effective natural language (NL) preference elicitation (PE) strategies that can efficiently learn a user’s top item preferences in cold start settings, ideally requiring only an arbitrary set of NL item descriptions. While the advent of large language models (LLMs) has introduced the technology to facilitate NL-PE conversations (Handa et al., 2024; Li et al., 2023) we conjecture that monolithic LLMs have limited abilities to strategically conduct active, multi-turn NL-PE dialogues about a set of arbitrary items. Specifically, we hypothesize that LLMs lack the sequential decision-theoretic reasoning to interactively generate queries that avoid over-exploitation or over-exploration of user-item preferences, thus risking over-focusing on already revealed item preferences or wastefully exploring preferences over low-value items. Further challenges faced by monolithic LLM NL-PE approaches include the need to jointly reason over large, potentially unseen sets of item descriptions, and the lack of control and interpretability in system behaviour even after prompt engineering or fine-tuning (Maynez et al., 2020).
@@
In contrast, conventional PE algorithms (Li et al., 2021; Zhang et al., 2020; Lei et al., 2020; Lee et al., 2019; Zhao et al., 2013), including Bayesian optimization methods (Guo and Sanner, 2010; Yang et al., 2021; Christakopoulou et al., 2016; Vendrov et al., 2019; Boutilier, 2002), establish formal decision-theoretic policies such as Thompson Sampling (TS) and Upper Confidence Bound (UCB) (Karamanolakis et al., 2018) to balance exploration and exploitation with the goal of quickly identifying the user’s most preferred items. However, these techniques typically assume a user can express preferences via direct item ratings or comparisons – an unrealistic expectation when users are unfamiliar with most items (Biyik et al., 2023). While recent work has extended Bayesian PE to a fixed set of template-based queries over pre-defined keyphrases (Yang et al., 2021), no existing work extends Bayesian methodologies to generative NL-PE over a set of generic NL item descriptions.
@@
  • We introduce the first Bayesian optimization formalization of NL-PE for arbitrary NL dialogue over a generic set of NL item descriptions – establishing a framework for research on augmenting LLMs with decision-theoretic reasoning.
@@
  • We numerically evaluate PEBOL against monolithic GPT-3.5 NL-PE methods via controlled NL-PE dialogue experiments over multiple NL item datasets and levels of user noise.
  • We observe that PEBOL achieves up to 131% improvement in MAP@10 after 10 turns compared to GPT-3.5, despite using a much smaller 400M NLI model for preference inference.
@@ 2.2. Preference Elicitation
PE has witnessed decades of research, and includes approaches based on Bayesian optimization (e.g., (Eric et al., 2007; Guo et al., 2010; Brochu et al., 2010; González et al., 2017; Khajah et al., 2016)), Bandits (e.g., (Li et al., 2010, 2011; Zhao et al., 2013; Christakopoulou et al., 2016)), constrained optimization (Rossi and Sperduti, 2004), and POMDPs (Boutilier, 2002). In the standard PE setting, a user is assumed to have some hidden utilities $\mathbf{u}=[u_{1},...,u_{N}]$ over a set $\mathcal{I}$ of $N$ items, where item $i$ is preferred to item $j$ if $u_{i}>u_{j}$. The goal of PE is typically to search for an item $i^{*}\in\arg\max_{i}u_{i}$ that maximizes user utility in a minimal number of PE queries, which most often ask a user to express item preferences as item ratings (e.g., (Li et al., 2010, 2011; Brochu et al., 2010; Zhao et al., 2013; Christakopoulou et al., 2016)) or relative preferences between item pairs or sets (e.g., (Boutilier, 2002; Guo and Sanner, 2010; Handa et al., 2024; Eric et al., 2007; González et al., 2017; Vendrov et al., 2019)). An alternative form of PE asks users to express preferences over predefined item features, also through rating- or comparison-based queries (Li et al., 2021; Zhang et al., 2020; Lei et al., 2020). Central to the above PE methods are query selection strategies that balance the exploration and exploitation of user preferences, with TS and UCB algorithms (cf. Sec. 4.2) often exhibiting strong performance (Zhao et al., 2013; Christakopoulou et al., 2016; Yang et al., 2021; Li et al., 2021; Zhang et al., 2020). However, none of these methods are able to interact with users through NL dialogue.
@@ 2.5. Natural Language Inference
Binary Natural Language Inference (NLI) (Yin et al., 2019) models predict the likelihood that one span of text called a premise is entailed by (i.e., can be inferred from) a second span called the hypothesis. For example, an effective NLI model should predict a high likelihood that the premise “I want to watch Iron Man” entails the hypothesis “I want to watch a superhero movie”. As illustrated by this example, the hypothesis typically must be more general than the premise. NLI models are trained by fine-tuning encoder-only LLMs on NLI datasets (Williams et al., 2017; Dagan et al., 2005; Thorne et al., 2018), which typically consist of short text spans for the premise and hypothesis – thus enabling relatively efficient performance on similar tasks with a fairly small number LLM parameters.
@@ 3. Problem Definition
To formulate NL-PE as a Bayesian optimization problem, we place a prior belief on the user’s utilities $p(\mathbf{u}|\mathbf{x})$, conditioned on item descriptions since they are available before the dialogue begins. We then assume an observation model that gives the likelihood $p(\mathbf{r}^{t}|\mathbf{x},\mathbf{u},\mathbf{q}^{t}$), letting us define the posterior utility belief as
@@ 4. Methodology
Figure 4. MAP@10 for MonoLLM and PEBOL-P with uncertainty-informed policies (UCB, TS, ER) – all methods show preference learning over time. MonoLLM is outperformed by PEBOL on Yelp and MovieLens, but remains competitive with PEBOL up to the fourth turn on Recipe-MPR until demonstrating a sharp performance drop that takes several turns to recover from.
@@ 4.1.1. Prior Beliefs
Before any dialogue, PEBOL establishes a prior belief $p(\mathbf{u}|\mathbf{x})$ on user-item utilities. We assume item utilities are independent so that
(6) $p(\mathbf{u}|\mathbf{x})=\prod_{i=1}^{N}p(u_{i}|x_{i}),$
@@
(7) $p(u_{i}|x_{i})=\text{Beta}(\alpha_{i}^{0},\beta_{i}^{0}).$
@@ 4.1.4. Natural Language Inference and Probabilistic Posterior Update
To deal with the fact that NLI models actually return an entailment probability, our probabilistic observation variant, PEBOL-P leverages the entailment probability that $r_{i}^{t}=1$ (cf. Sec 4.3), which we denote as $w_{i}^{t}\in[0,1]$. We then simply relax our previous binary observations $r_{i}^{t}$ to continuous $w_{i}^{t}$, thus leading to soft posterior updates $\alpha_{i}^{t}=\alpha_{i}^{t-1}+w_{i}^{t}$, $\beta_{i}^{t}=\beta_{i}^{t-1}+(1-w_{i}^{t})$.
@@
Figure 5. MAP@10 for PEBOL-P with various context acquisition policies.
@@ 4.3.2. Inferring Item Ratings from NL Preferences
Given a NL preference $\rho^{t}$, PEBOL must infer whether the user would like an item described by $x_{i}$. Specifically, PEBOL acquires ratings $\mathbf{w}^{t}=[w^{t}_{1},...,w^{t}_{N}]$ (cf. Sec. 4.1.4) by using NLI to predict whether an item description $x_{i}$ entails (i.e., implies) the preference $\rho^{t}$. For example, we expect that an NLI model would predict that $x_{i}=$The Lion King” entails $\rho^{t}=$“animated” while $x_{j}=$“Titanic” does not, inferring that a user who expressed preference $\rho^{t}$ would like item $i$ but not $j$. We use an NLI model $g_{\theta}(x_{i},\rho^{t})$ to predict the probability $w_{i}^{t}$ that $x_{i}$ entails $\rho^{t}$, and return $r_{i}^{t}=\lfloor w_{i}^{t}\rceil$ in the case of binary observations (PEBOL-B) and $w_{i}^{t}$ in the case of probabilistic observations (PEBOL-P).
@@ 4.4. The Complete PEBOL System
Figure 6. MAP@10 for PEBOL using binary vs. probabilistic entailment scores. PEBOL-P with the best policy (TS on Yelp and MovieLens, UCB on Recipe-MPR) generally outperforms PEBOL-B.
@@ 5. Experimental Methods
We numerically evaluate our PEBOL variations through controlled NL-PE dialogue experiments across multiple datasets and response noise levels – comparing against a monolithic GPT-3.5-turbo NL-PE system, which we refer to as MonoLLM. We do not compare against ConvRec methods (Li et al., 2018; Chen et al., 2019; Wang et al., 2022; Yang et al., 2022) because they are not cold-start systems, requiring observed user-item interactions data to drive their recommendation modules. We also do not base our experiments on ConvRec datasets such as ReDIAL (Li et al., 2018), since they are made up of pre-recorded conversation histories and cannot be used to evaluate active, cold-start NL-PE systems.
@@ 5.1. MonoLLM Baseline
A major challenge of using MonoLLM for NL-PE is that item descriptions $\mathbf{x}$ either need to be internalized through training or be provided in the context window (cf. Sec. 4). Since we focus on fully cold-start settings, we test the latter approach as a baseline. Specifically, we prompt the LLM to generate new queries to elicit preferences given the full conversation history $\mathcal{H}^{t}$ and $\mathbf{x}$ (all prompts are shown in the Supplementary Materials). We evaluate recommendation performance after each turn by using another prompt to recommend a list of ten item names from $\mathbf{x}$ given $\mathcal{H}^{t}$. Due to context window limits, this MonoLLM approach is only feasible for small item sets with short item descriptions; thus, we have to limit $|\mathcal{I}|$ to 100 for fair comparison to the MonoLLM baseline.
@@ 5.2. Simulation Details
We test PEBOL and MonoLLM through NL-PE dialogues with LLM-simulated users whose item preferences are hidden from the system, evaluating recommendation performance over 10 turns of dialogue, and using GPT-3.5-turbo-0613 for all text generation.
@@ 5.2.2. PEBOL Query Generation
Given an item description $x_{i}$ and previously generated aspects $(a^{1},...,a^{t-1})$, an LLM is prompted to generate an aspect $a^{t}$ describing the item $i$ that is no more than 3 words long. The LLM is then prompted again to generate a “yes-or-no” query asking if a user prefers $a^{t}$.

### v2

@@ Abstract.
Designing preference elicitation (PE) methodologies that can quickly ascertain a user’s top item preferences in a cold-start setting is a key challenge for building effective and personalized conversational recommendation (ConvRec) systems. While large language models (LLMs) enable fully natural language (NL) PE dialogues, we hypothesize that monolithic LLM NL-PE approaches lack the multi-turn, decision-theoretic reasoning required to effectively balance the exploration and exploitation of user preferences towards an arbitrary item set. In contrast, traditional Bayesian optimization PE methods define theoretically optimal PE strategies, but cannot generate arbitrary NL queries or reason over content in NL item descriptions – requiring users to express preferences via ratings or comparisons of unfamiliar items. To overcome the limitations of both approaches, we formulate NL-PE in a Bayesian Optimization (BO) framework that seeks to actively elicit NL feedback to identify the best recommendation. Key challenges in generalizing BO to deal with natural language feedback include determining: (a) how to leverage LLMs to model the likelihood of NL preference feedback as a function of item utilities, and (b) how to design an acquisition function for NL BO that can elicit preferences in the infinite space of language. We demonstrate our framework in a novel NL-PE algorithm, PEBOL, which uses: 1) Natural Language Inference (NLI) between user preference utterances and NL item descriptions to maintain Bayesian preference beliefs, and 2) BO strategies such as Thompson Sampling (TS) and Upper Confidence Bound (UCB) to guide LLM query generation. We numerically evaluate our methods in controlled simulations, finding that after 10 turns of dialogue, PEBOL can achieve an MRR@10 of up to 0.27 compared to the best monolithic LLM baseline’s MRR@10 of 0.17, despite relying on earlier and smaller LLMs. [^1] Our code is publically available at https://github.com/D3Mlab/llm-pe.
@@
journalyear: 2024copyright: acmlicensedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688142isbn: 979-8-4007-0505-2/24/1 ccs: Information systems Recommender systemsccs: Information systems Personalizationccs: Information systems Language models
@@ 1. Introduction
Personalized conversational recommendation (ConvRec) systems require effective natural language (NL) preference elicitation (PE) strategies that can efficiently learn a user’s top item preferences in cold start settings, ideally requiring only an arbitrary set of NL item descriptions. While the advent of large language models (LLMs) has introduced the technology to facilitate NL-PE conversations (Handa et al., 2024; Li et al., 2023) we conjecture that monolithic LLMs have limited abilities to strategically conduct active, multi-turn NL-PE dialogues about a set of arbitrary items. Specifically, we hypothesize that LLMs lack the multi-turn decision-theoretic reasoning to interactively generate queries that avoid over-exploitation or over-exploration of user-item preferences, thus risking over-focusing on already revealed item preferences or wastefully exploring preferences over low-value items. Further challenges faced by monolithic LLM NL-PE approaches include the need to jointly reason over large, potentially unseen sets of item descriptions, and the lack of control and interpretability in system behaviour even after prompt engineering or fine-tuning (Maynez et al., 2020).
@@
In contrast, conventional PE algorithms (Li et al., 2021; Zhang et al., 2020; Lei et al., 2020; Lee et al., 2019; Zhao et al., 2013), including Bayesian optimization methods (Guo and Sanner, 2010; Yang et al., 2021; Christakopoulou et al., 2016; Vendrov et al., 2020; Boutilier, 2002), establish formal decision-theoretic policies such as Thompson Sampling (TS) and Upper Confidence Bound (UCB) (Karamanolakis et al., 2018) to balance exploration and exploitation with the goal of quickly identifying the user’s most preferred items. However, these techniques typically assume a user can express preferences via direct item ratings or comparisons – an unrealistic expectation when users are unfamiliar with most items (Biyik et al., 2023). While recent work has extended Bayesian PE to a fixed set of template-based queries over pre-defined keyphrases (Yang et al., 2021), no existing work extends Bayesian methodologies to generative NL-PE over a set of generic NL item descriptions.
@@
  • We introduce the first Bayesian optimization formalization of NL-PE for arbitrary NL dialogue over a generic set of NL item descriptions – establishing a new framework for research on steering LLMs with decision-theoretic reasoning.
@@
  • We numerically evaluate PEBOL against monolithic GPT-3.5 and Gemini-Pro NL-PE methods via controlled NL-PE dialogue experiments over multiple NL item datasets and levels of user noise.
  • We observe that after 10 turns of dialogue, PEBOL can achieve a mean MRR@10 of up to 0.27 compared to the best monolithic LLM baseline’s MRR@10 of 0.17, despite relying on earlier and smaller LLMs.
@@ 2.2. Preference Elicitation
PE has witnessed decades of research, and includes approaches based on Bayesian optimization (e.g., (Eric et al., 2007; Guo et al., 2010; Brochu et al., 2010; González et al., 2017; Khajah et al., 2016)), Bandits (e.g., (Li et al., 2010, 2011; Zhao et al., 2013; Christakopoulou et al., 2016)), constrained optimization (Rossi and Sperduti, 2004), and POMDPs (Boutilier, 2002). In the standard PE setting, a user is assumed to have some hidden utilities $\mathbf{u}=[u_{1},...,u_{N}]$ over a set $\mathcal{I}$ of $N$ items, where item $i$ is preferred to item $j$ if $u_{i}>u_{j}$. The goal of PE is typically to search for an item $i^{*}\in\arg\max_{i}u_{i}$ that maximizes user utility in a minimal number of PE queries, which most often ask a user to express item preferences as item ratings (e.g., (Li et al., 2010, 2011; Brochu et al., 2010; Zhao et al., 2013; Christakopoulou et al., 2016)) or relative preferences between item pairs or sets (e.g., (Boutilier, 2002; Guo and Sanner, 2010; Handa et al., 2024; Eric et al., 2007; González et al., 2017; Vendrov et al., 2020)). An alternative form of PE asks users to express preferences over predefined item features, also through rating- or comparison-based queries (Li et al., 2021; Zhang et al., 2020; Lei et al., 2020). Central to the above PE methods are query selection strategies that balance the exploration and exploitation of user preferences, with TS and UCB algorithms (cf. Sec. 4.2) often exhibiting strong performance (Zhao et al., 2013; Christakopoulou et al., 2016; Yang et al., 2021; Li et al., 2021; Zhang et al., 2020). However, none of these methods are able to interact with users through NL dialogue or reason about NL item descriptions.
@@ 2.5. Natural Language Inference
Binary Natural Language Inference (NLI) (Yin et al., 2019) models predict the likelihood that one span of text called a premise is entailed by (i.e., can be inferred from) a second span called the hypothesis. For example, an effective NLI model should predict a high likelihood that the premise “I want to watch Iron Man” entails the hypothesis “I want to watch a superhero movie”. As illustrated by this example, the hypothesis typically must be more general than the premise. NLI models are trained by fine-tuning encoder-only LLMs on NLI datasets (Williams et al., 2018; Dagan et al., 2005; Thorne et al., 2018), which typically consist of short text spans for the premise and hypothesis – thus enabling relatively efficient performance on similar tasks with a fairly small number LLM parameters.
@@ 3. Problem Definition
To formulate NL-PE as a Bayesian optimization problem, we place a prior belief on the user’s utilities $p(\mathbf{u}|\mathbf{x})$, potentially conditioned on item descriptions since they are available before the dialogue begins. We then assume an observation model that gives the likelihood $p(\mathbf{r}^{t}|\mathbf{x},\mathbf{u},\mathbf{q}^{t}$), letting us define the posterior utility belief as
@@ 4. Methodology
Figure 4. MRR@10 for MonoLLM and PEBOL-P with uncertainty-informed policies (UCB, TS, ER). All methods show preference learning over time and MonoLLM is generally outperformed by PEBOL.
@@ 4.1.1. Prior Beliefs
Before any dialogue, PEBOL establishes an uninformed prior belief $p(\mathbf{u})$ on user-item utilities. We assume item utilities are independent so that
(6) $p(\mathbf{u})=\prod_{i=1}^{N}p(u_{i}),$
@@
(7) $p(u_{i})=\text{Beta}(\alpha_{i}^{0},\beta_{i}^{0}).$
@@ 4.1.4. Natural Language Inference and Probabilistic Posterior Update
To deal with the fact that NLI models actually return an entailment probability, our probabilistic observation variant, PEBOL-P leverages the probability that item description $x_{i}$ entails $q_{t}$, which we denote as $w_{i}^{t}$. We provide a full graphical model and derivation of the Bayesian posterior update given this entailment probability in the Supplementary Material, but note that we can summarize the final result as a relaxed version of the binary posterior update of (13) that replaces the binary observation $r_{i}\in\{0,1\}$ with the entailment probability $w_{i}^{t}\in[0,1]$, i.e., $\alpha_{i}^{t}=\alpha_{i}^{t-1}+w_{i}^{t}$, $\beta_{i}^{t}=\beta_{i}^{t-1}+(1-w_{i}^{t})$.
@@
Figure 5. MRR@10 for PEBOL-P with various context acquisition policies.
@@ 4.3.2. Inferring Item Ratings from NL Preferences
Given a NL preference $\rho^{t}$, PEBOL must infer whether the user would like an item described by $x_{i}$. Specifically, PEBOL acquires ratings $\mathbf{w}^{t}=[w^{t}_{1},...,w^{t}_{N}]$ (cf. Sec. 4.1.4) by using NLI to predict whether an item description $x_{i}$ entails (i.e., implies) the preference $\rho^{t}$. For example, we expect that an NLI model would predict that $x_{i}=$The Lion King” entails $\rho^{t}=$“animated” while $x_{j}=$“Titanic” does not, inferring that a user who expressed preference $\rho^{t}$ would like item $i$ but not $j$. We use an NLI model $P_{\omega}(x_{i},\rho^{t})$ to predict the probability $w_{i}^{t}$ that $x_{i}$ entails $\rho^{t}$, and return $r_{i}^{t}=\lfloor w_{i}^{t}\rceil$ in the case of binary observations (PEBOL-B) and $w_{i}^{t}$ in the case of probabilistic observations (PEBOL-P).
@@ 4.4. The Complete PEBOL System
Figure 6. MRR@10 for PEBOL using binary vs. probabilistic entailment scores. PEBOL-P with the best policy (TS on Yelp and MovieLens, UCB on Recipe-MPR) generally outperforms PEBOL-B.
@@ 5. Experimental Methods
We numerically evaluate our PEBOL variations through controlled NL-PE dialogue experiments across multiple datasets and response noise levels – comparing against two monolithic LLM (MonoLLM) baselines. Specifically, these baselines directly use GPT-3.5-turbo-0613 (GPT MonoLLM) or Gemini-Pro (Gemini MonoLLM) as the NL-PE system, as described in Section 5.1. We do not compare against ConvRec methods (Li et al., 2018; Chen et al., 2019; Wang et al., 2022; Yang et al., 2022) because they are not cold-start systems, requiring observed user-item interactions data to drive their recommendation modules. We also do not base our experiments on ConvRec datasets such as ReDIAL (Li et al., 2018), since they are made up of pre-recorded conversation histories and cannot be used to evaluate active, cold-start NL-PE systems.
@@ 5.1. MonoLLM Baseline
A major challenge of using MonoLLM for NL-PE is that item descriptions $\mathbf{x}$ either need to be internalized through training or be provided in the context window (cf. Sec. 4) – since we focus on fully cold-start settings, we test the latter approach as a baseline. In each turn, given the full conversation history $\mathcal{H}^{t}$ and $\mathbf{x}$, we prompt the MonoLLM to generate a new query to elicit user preferences – all prompts are shown in the Supplementary Materials. We evaluate recommendation performance after each turn by using another prompt to recommend a list of ten item names from $\mathbf{x}$ given $\mathcal{H}^{t}$. Due to context window limits, this MonoLLM approach is only feasible for small item sets with short item descriptions; thus, we have to limit $|\mathcal{I}|$ to 100 for fair comparison to the MonoLLM baseline.
@@ 5.2. Simulation Details
We test PEBOL and MonoLLM through NL-PE dialogues with LLM-simulated users, where the simulated users’ item preferences remain hidden from the system. We evaluate recommendation performance over 10 turns of dialogue.
@@ 5.2.1. User Simulation
5.2.2. Evaluating Recommendations
We evaluate the top-10 recommendations in each turn using the Mean Reciprocal Rank (MRR@10) of the preferred item, which is equivalent to MAP@10 for the case of a single preferred item.
@@ 5.2.3. PEBOL Query Generation
In turn $t$, given an item description $x_{i}$ and previously generated aspects $(a^{1},...,a^{t-1})$, an LLM (GPT-3.5-turbo-0613) [^5] Experiments with newer LLMs such as Gemini or GPT4 for PEBOL query generation are left for future work due to the API time and cost requirements needed to simulate the many variants of PEBOL reported in Section 6. We do, however, compare PEBOL against a Gemini-MonoLLM baseline. is prompted to generate an aspect $a^{t}$ describing the item $i$ that is no more than 3 words long. The LLM is then prompted again to generate a “yes-or-no” query asking if a user prefers $a^{t}$.